OpenStudy (anonymous):

Another question for midterm exam studying. derivative of ln(x^(2)e^(x))

hartnn (hartnn):

simplify ln(x^(2)e^(x)) using log properties, can you ?

hartnn (hartnn):

ln (AB) = ln A+ln b

OpenStudy (anonymous):

so ln(x^2)+ln(e^x)

hartnn (hartnn):

yes, now use ln m^n = n ln m and ln e=1

OpenStudy (anonymous):

so 2ln(x)+1

hartnn (hartnn):

ln e^x = x ln e = x so, 2 ln x+x

hartnn (hartnn):

can you differentiate that ^ ?

OpenStudy (anonymous):

doesnt ln(x)'=1/x? So the answer would be 2/x +1 right?

hartnn (hartnn):

thats correct. that is also correct :)

OpenStudy (anonymous):

Awesome thank you I'll be posting probably about 2 more questions that im stuck on. Thanks again for your help

hartnn (hartnn):

welcome ^_^